I am trying to learn the elimination method. Please tell me what I did wrong.
5x+3y=-13
7x-2y=17
10x+6y=-26
21x-6y=51
31x=25
x= 25/31
5x+3y=-13
5(25/31)+3y=-13
125/31+3y=-13
3y=-13-125/31
that is where I get stuck. Can you help?

Answer by Alan3354(69443)

5x+3y=-13
7x-2y=17
-------------
10x+6y=-26
21x-6y=51
----------
31x=25
x= 25/31
----------
5x+3y=-13
5(25/31)+3y=-13
125/31+3y=-13
3y=-13 - (125/31)
3y = (-403 - 125)/13 = -528/13
y = -176/39
You were almost there.
that is where I get stuck. Can you help?